OpenStudy (anonymous):

Need help figuring this out i dont want the answer up front i want someone to help me work it out Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

well slope would be 3 plug into point-slope form y+2=3(x-4) solve for y

OpenStudy (the_fizicx99):

@mathman806 wouldn't the slope be -3 because you subtract 3x for y = mx+b

OpenStudy (anonymous):

well you subtract 3x then you have -y=-3x-3 and so you divide by -1 to get the y by itself y=3x+3 @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Oh -y/-1 = -3x/-1 -3/-1 :D nvm *hides back in closet*

OpenStudy (anonymous):

wait where do you get -1? the invis -1 in front of -y? so -y is = to -1y?

OpenStudy (the_fizicx99):

To cancel out -y

OpenStudy (the_fizicx99):

-y is the same thing as -1y

OpenStudy (anonymous):

ok so then i ended up with -y= -3x-3 and then i divide -1 to get y=-3x+3 ?

OpenStudy (anonymous):

the -3x will become 3x

OpenStudy (the_fizicx99):

No, y = 3x + 3. Look at what mathman posted above

OpenStudy (anonymous):

then where does the +3 come from?

OpenStudy (anonymous):

-3/-1=3 all terms get divided by -1

OpenStudy (anonymous):

OOOHHHHH OK haha i feel dum. so then what do i do with y=3x+3

OpenStudy (anonymous):

using the 3 from the 3x and the point put it into point-slope form y+2=3(x-4) solve for y

OpenStudy (anonymous):

how'd it go from y=3x+3 to y+2=3(x-4) ? o.O

OpenStudy (anonymous):

yell point-slope form is y-y1=m(x-x1) y1 is the y-value x1 is the x-value m is the slope from the first equation

OpenStudy (anonymous):

y+2=3(x-4) <-- why does the formula say y-y1 but the problem says y+2

OpenStudy (anonymous):

2 being y1

OpenStudy (anonymous):

@mathman806

OpenStudy (anonymous):

anyone? what do i do with this ? y+2=3(x-4)

OpenStudy (anonymous):

take the negivte reciprocal of the slope

OpenStudy (anonymous):

cause the 2 is a negative 2 and 2 negatives makes a positive -(-2)=2 @zekewars